Find the particular solution of:

1.) x^2 y' -2xy=x^4 +3; where y = 2 and x = 1


1
Expert's answer
2022-03-08T02:20:02-0500
"x^2y'-2xy=x^4+3"

"y'-\dfrac{2}{x}=x^2+\dfrac{3}{x^2}"

"I.F.=e^{\int(-2/x)dx}=\dfrac{1}{x^2}"

"\dfrac{1}{x^2}(y'-\dfrac{2}{x})=\dfrac{1}{x^2}(x^2+\dfrac{3}{x^2})"


"d(\dfrac{y}{x^2})=(1+\dfrac{3}{x^4})dx"

Integrate


"\int d(\dfrac{y}{x^2})=\int(1+\dfrac{3}{x^4})dx"

"\dfrac{y}{x^2}=x-\dfrac{1}{x^3}+C"

"y=x^3-\dfrac{1}{x}+Cx^2"

"y = 2,x = 1"


"2=1^3-\dfrac{1}{1}+C(1)^2"

"C=2"

The particular solution is


"y=x^3-\dfrac{1}{x}+2x^2"